If it takes Roger 60 minutes to paint a wall. It takes Stephanie 40 minutes to paint the same wall. How long will it take Roger and Stephanie to paint the wall if the work together?
Answer by fractalier(6550)

These work problems are set up the same way...
t/A + t/B = 1
where t is the time worked together, and A and B are how lojng it takes themn by themselves and the 1 represents one job...so plugging in we get
t/40 + t/60 = 1 multiply by 120
3t + 2t = 120
5t = 120
t = 24 min